# Quantum-HPC Framework with multi-GPU-Enabled Hybrid Quantum-Classical Workflow:

### 09 Mar 2024 v1

@@ Abstract
Quantum Machine Learning, Distributed Quantum Computing, Quantum Software, Quantum Simulation
@@ I Introduction
In this study, we endeavor to develop a framework that is conducive to high-performance computing (HPC) within the quantum domain [19], specifically designed to enable distributed quantum computing through a hybrid quantum-classical workflow. This framework incorporates multi-GPU acceleration (via the cuQuantum SDK[20]) to support quantum simulators (which serve as counterparts to noiseless QPUs, if QPUs are not available). Additionally, we integrate the latest Pennylane Lightning plugins[21]. This component leverages CUDA-aware MPI (Message Passing Interface) boosted by NVLink’s 600 GB/s bidirectional bandwidth for optimized GPU-to-GPU communication, enabling rapid distributed simulation tasks[22].
@@ II-B Hamiltonians for Solving Phase Transition
II-C Towrad Distributed Quantum Computing
Fig. 4 delineates an integrated architecture designed to leverage distributed quantum resources for advanced computational tasks. The schematic begins with a classical scheduling layer, where a CPU disaggregates a complex problem into sub-problems. These are then converted into quantum states through GPUs in an encoding layer, facilitating the classical-to-quantum transition. The middle of the system can lie in the distributed-VQE layer[34, 24], where QPUs (linked by quantum internet and augmented by quantum repeaters) execute quantum algorithms. Finally, a hybrid quantum-classical neural network analyzes the quantum states, yielding a scalar output via a quantum classifier. This streamlined, multi-layered framework illustrates a scalable approach for executing computationally intensive tasks across distributed quantum systems, aiming to unlock new potentials in near-term quantum information processing.

### 18 Mar 2024 v2

# Multi-GPU-Enabled Hybrid Quantum-Classical Workflow in Quantum-HPC Middleware:

@@ Abstract
Quantum Machine Learning, Distributed Quantum Computing, Quantum Software, Quantum Simulation, cuQuantum SDK, Tensor Network
@@ I Introduction
In this study, we endeavor to develop a scheme that is conducive to high-performance computing (HPC) within the quantum domain [19], specifically designed to enable distributed quantum computing through a hybrid quantum-classical workflow. This framework incorporates multi-GPU acceleration (via the cuQuantum SDK[20]) to support quantum simulators (which serve as counterparts to noiseless QPUs, if QPUs are not available). Additionally, we integrate the latest Pennylane Lightning plugins[21]. This component leverages CUDA-aware MPI (Message Passing Interface) boosted by NVLink’s 600 GB/s bidirectional bandwidth for optimized GPU-to-GPU communication, enabling rapid distributed simulation tasks[22]. This middleware scheme offers a synergy between classical GPU acceleration and quantum computing acceleration[23, 19]. The relevant conceptual architectures for Quantum-HPC middleware are discussed by Saurabh et. al[19].
@@ II-B Hamiltonians for Solving Phase Transition
II-C Towrad Distributed Quantum Computing: a Conceptual Architecture for a Distributed-Quantum-HPC Middleware
To develop a conceptual architecture for distributed quantum computing with a QCQ workflow, we expand upon the concept based on the work of Saurabh et al[19]. This approach enables a deeper integration of quantum computing processes within distributed computing frameworks. Based on Fig. 4 delineates an integrated architecture designed to leverage distributed quantum resources for advanced computational tasks. The schematic begins with a classical scheduling layer, where a CPU disaggregates a complex problem into sub-problems. These are then converted into quantum states through GPUs in an encoding layer, facilitating the classical-to-quantum transition. The middle of the system can lie in the distributed-VQE layer[35, 25], where QPUs (linked by quantum internet and augmented by quantum repeaters) execute quantum algorithms. Finally, a hybrid quantum-classical neural network analyzes the quantum states, yielding a scalar output via a quantum classifier. This streamlined, multi-layered framework illustrates a scalable approach for executing computationally intensive tasks across distributed quantum systems, aiming to unlock new potentials in near-term quantum information processing.
@@ IV Conclusion
V Future work
Our upcoming research aims to enhance the QCQ architecture by harnessing the CUDA Quantum platform[44], delivering substantial speedup and flexibility for quantum-classical computing workflows through distributed quantum simulation and seamless integration with machine learning frameworks.